### v2

@@ Abstract
In a graph, a (perfect) matching cut is an edge cut that is a (perfect) matching. matching cut (mc), respectively, perfect matching cut (pmc), is the problem of deciding whether a given graph has a matching cut, respectively, a perfect matching cut. The disconnected perfect matching problem (dpm) is to decide if a graph has a perfect matching that contains a matching cut. Solving an open problem posed in [Lucke, Paulusma, Ries (ISAAC 2022, Algorithmica 2023) & Feghali, Lucke, Paulusma, Ries (arXiv:2212.12317)], we show that pmc is $\mathsf{NP}$-complete in graphs without induced $14$-vertex path $P_{14}$. Our reduction also works simultaneously for mc and dpm, improving the previous hardness results of mc on $P_{15}$-free graphs and of dpm on $P_{19}$-free graphs to $P_{14}$-free graphs for both problems. Actually, we prove a slightly stronger result: within $P_{14}$-free $8$-chordal graphs (graphs without chordless cycles of length at least $9$), it is hard to distinguish between those without matching cuts (respectively, perfect matching cuts, disconnected perfect matchings) and those in which every matching cut is a perfect matching cut. Moreover, assuming the Exponential Time Hypothesis, none of these problems can be solved in $2^{o(n)}$ time for $n$-vertex $P_{14}$-free $8$-chordal graphs.
On the positive side, we show that, as for mc [Moshi (JGT 1989)], dpm and pmc are polynomially solvable when restricted to $4$-chordal graphs. Together with the negative results, this partly answers an open question on the complexity of pmc in $k$-chordal graphs asked in [Le, Telle (TCS 2022) & Lucke, Paulusma, Ries (MFCS 2023)].
@@ 1 Introduction and results
In a graph $G=(V,E)$, a cut is a partition $V=X\cup Y$ of the vertex set into disjoint, non-empty sets $X$ and $Y$. The set of all edges in $G$ having an endvertex in $X$ and the other endvertex in $Y$, written $E(X,Y)$, is called the edge cut of the cut $(X,Y)$. A matching cut is an edge cut that is a (possibly empty) matching. Another way to define matching cuts is as follows; see Chvatal84 ; Graham70 : a cut $(X,Y)$ is a matching cut if and only if each vertex in $X$ has at most one neighbor in $Y$ and each vertex in $Y$ has at most one neighbor in $X$. The classical $\mathsf{NP}$-complete problem matching cut (mcChvatal84 asks if a given graph admits a matching cut.
An interesting special case, where the edge cut $E(X,Y)$ is a perfect matching of $G$, was considered in HeggernesT98 . Such a matching cut is called a perfect mathing cut and the perfect matching cut (pmc) problem asks whether a given graph admits a perfect matching cut. It was shown in HeggernesT98 that this special case pmc of mc remains $\mathsf{NP}$-complete.
A notion related to matching cut is disconnected perfect matching which has been considered recently in BouquetP21 : a disconnected perfect matching is a perfect matching that contains a matching cut. Observe that any perfect matching cut is a disconnected perfect matching but not the converse. Fig. 1 provides some small examples for matching cuts, perfect matching cuts and disconnected perfect matchings.
(a)(b)(c)(d)
@@
The related problem to mc and pmc, disconnected perfect matching (dpm), asks if a given graph has a disconnected perfect matching; equivalently: if a given graph has a matching cut that is extendable to a perfect matching. It was shown in BouquetP21 that dpm is $\mathsf{NP}$-complete. All these three problems have received much attention lately; see, e.g.,  BonnetCD23 ; BouquetP21 ; ChenHLLP21 ; FeghaliLPR22 ; FeghaliLPR23 ; GolovachKKL22 ; LeT22 ; LuckePR23 ; LuckePR23J for recent results. In particular, we refer to LuckePR23J for a short survey on known applications of matching cuts in other research areas.
In this paper, we focus on the complexity and algorithms of these three problems restricted to graphs without long induced paths and cycles. In Feghali23 , Feghali initiated the study of mc for graphs without a fixed path as an induced subgraph. He showed that mc is solvable in polynomial time for $P_{5}$-free graphs, but that there exists an integer $t>0$ for which it is $\mathsf{NP}$-complete for $P_{t}$-free graphs; indeed, $t=27$ as pointed out in LuckePR23J . The current best known hardness results for mc and dpm in graphs without long induced paths are:
Theorem 1 (LuckePR23J )
@@
Prior to the present paper, no similar hardness result for pmc was known. Indeed, it was asked in FeghaliLPR22 ; LuckePR22isaac ; LuckePR23J , whether there is an integer $t$ such that pmc is $\mathsf{NP}$-complete in $P_{t}$-free graphs. Polynomial-time algorithms exist for mc, pmc and dpm in $P_{6}$-free graphs LuckePR22isaac ; LuckePR22arxiv ; LuckePR23 .
@@
Previously, no similar polynomial-time results for dpm and pmc in $4$-chordal were known. Indeed, the computational complexity of pmc in $k$-chordal graphs, $k\geq 4$, has been stated as an open problem in LeT22 ; LuckePR23 .
Our contributions.
We prove that pmc is $\mathsf{NP}$-complete in graphs without induced path $P_{14}$, solving the open problem posed in FeghaliLPR22 ; LuckePR22isaac ; LuckePR22arxiv ; LuckePR23J . For mc and dpm  we improve the hardness results in Theorem 1 in graphs without induced path $P_{15}$, respectively, $P_{19}$, to graphs without induced path $P_{14}$. It is remarkable that all these hardness results for three problems will be obtained simultaneously by only one reduction. Thus, our approach unifies and improves known approaches and results in the literature. Our hardness results can be stated in more details as follows.
@@
Our polynomial-time algorithm solving dpm in $4$-chordal graphs is based on simple, known forcing rules. We remark that the approach also works for mc, providing another proof for Theorem 2. In case of pmc in $4$-chordal graphs, we will give a non-trivial polynomial-time reduction to 2sat, based on the distance-level structure of $4$-chordal graphs admitting perfect matching cuts. Our positive and negative results in Theorems 3 and 4 on $k$-chordal graphs partly answer an open question stated in LeT22 ; LuckePR23 .
@@
$t$ $\leq 5$ $6$ $7$cdots $13$ $\geq 14$
mc $\mathsf{P}$ $\mathsf{P}$cdots ? $\mathsf{NPC}$()($\star$)( ⋆ )
pmc $\mathsf{P}$ $\mathsf{P}$cdots ? $\mathsf{NPC}$()($\star$)( ⋆ )
dpm $\mathsf{P}$ $\mathsf{P}$cdots ? $\mathsf{NPC}$()($\star$)( ⋆ )
$k$ $\leq 4$ $5$ $6$ $7$ $\geq 8$ mc $\mathsf{P}$ $\mathsf{NPC}$()($\star$)( ⋆ ) pmc $\mathsf{P}$()($\star$)( ⋆ ) ? ? ? $\mathsf{NPC}$()($\star$)( ⋆ ) dpm $\mathsf{P}$()($\star$)( ⋆ ) ? ? ? $\mathsf{NPC}$()($\star$)( ⋆ )
Table 1: The current status of the computational complexity of matching cut problems in $P_{t}$-free graphs (left hand-side) and in $k$-chordal graphs (right hand-side). $\mathsf{P}$, $\mathsf{NPC}$, and ‘?’ means polynomial, $\mathsf{NP}$-complete, and unknown, respectively. Results of this paper are marked with ()($\star$)( ⋆ ).
@@ 2 Preliminaries
The neighborhood of a vertex $v$ in a graph $G$ is denoted $N_{G}(v)$, or simply $N(v)$ if the context is clear. We write $N[v]$ for $N(v)\cup\{v\}$. For a set $\cal H$ of graphs, ${\cal H}$-free graphs are those in which no induced subgraph is isomorphic to a graph in $\cal H$. We denote by $P_{t}$ the $t$-vertex path with $t-1$ edges and by $C_{t}$ the $t$-vertex cycle with $t$ edges. The $3$-vertex cycle $C_{3}$ is also called a triangle. The union $G+H$ of two vertex-disjoint graphs $G$ and $H$ is the graph with vertex set $V(G)\cup V(H)$ and edge set $E(G)\cup E(H)$; we write $pG$ for the union of $p$ copies of $G$. For a subset $S\subseteq V(G)$, let $G[S]$ denote the subgraph of $G$ induced by $S$; $G-S$ stands for $G[V(G)\setminus S]$. By ‘$G$ contains an $H$’ we mean $G$ contains $H$ as an induced subgraph.
@@
If $G$ has a perfect matching cut $M=(X,Y)$ then $N[v]$ is not monochromatic for any vertex $v$. Indeed, every vertex $v\in X$ has exactly one neighbor in $Y$ which we sometime call the private neighbor of $v$ and denote $p(v)$, and vice versa, every vertex $w\in Y$ has its private neighbor $p(w)$ in $X$. Thus, for any vertex $v$, $p(p(v))=v$.
Algorithmic lower bounds in this paper are conditional, based on the Exponential Time Hypothesis (ETH) ImpagliazzoP01 . The ETH asserts that no algorithm can solve 3sat in subexponential time $2^{o(n)}$ for $n$-variable 3-cnf formulas: there exists a constant $\delta>0$ such that no algorithm can solve 3sat in $2^{\delta n}$ time for 3-cnf formulas over $n$ variables. As shown by the Sparsification Lemma in ImpagliazzoPZ01 , the hard cases of 3sat already consist of sparse formulas with $m=O(n)$ clauses. Hence, the ETH implies that 3sat cannot be solved in $2^{o(n+m)}$ time.
Recall that an instance for 1-in-3sat is a 3-cnf formula $\phi=C_{1}\land C_{2}\land\cdots\allowbreak\land C_{m}$ over $n$ variables, in which each clause $C_{j}$ consists of three distinct literals. The problem asks whether there is a truth assignment to the variables such that every clause in $\phi$ has exactly one true literal. We call such an assignment an 1-in-3 assignment. There is a polynomial reduction from 3sat to 1-in-3sat ((Moret98, , Theorem 7.2), see also Schaefer78 ), which transforms an instance for 3sat to an equivalent instance for 1-in-3sat  linear in the number of variables and clauses. Thus, assuming ETH, 1-in-3sat cannot be solved in $2^{o(n+m)}$ time on inputs with $n$ variables and $m$ clauses. We will need a restriction of 1-in-3sat, positive 1-in-3sat, in which each variable occurs positively. There is a well-known polynomial reduction from 1-in-3sat to positive 1-in-3sat, which transforms an instance for 1-in-3sat to an equivalent instance for positive 1-in-3sat, linear in the number of variables and clauses. (See also an exercise in (Moret98, , Exercise 7.1).) Hence, we obtain: assuming ETH, positive 1-in-3sat cannot be solved in $2^{o(n+m)}$ time for inputs with $n$ variables and $m$ clauses.
3 Proof of Theorem 3
Recall that a perfect matching cut is in particular a matching cut, as well as a disconnected perfect matching. This observation leads to the following promise versions of mc, pmc and dpm. (We refer to Goldreich06a for background on promise problems.)
promise-pmc mc Instance: A graph $G$ that either has no matching cut, or every matching cut is a perfect matching cut. Question: Does $G$ have a matching cut ?
promise-pmc pmc Instance: A graph $G$ that either has no perfect matching cut, or every matching cut is a perfect matching cut. Question: Does $G$ have a perfect matching cut ?
promise-pmc dpm Instance: A graph $G$ that either has no disconnected perfect matching, or every matching cut is a perfect matching cut. Question: Does $G$ have a disconnected perfect matching ?
@@ 3.1 The reduction
Let $\phi$ be a 3-cnf formula with $m$ clauses $C_{j}$, $1\leq j\leq m$, and $n$ variables $x_{i}$, $1\leq i\leq n$, in which each clause $C_{j}$ consists of three distinct variables. We will construct a $\{3P_{6},2P_{7},P_{14}\}$-free $8$-chordal graph $G$ such that $G$ has a perfect matching cut if and only if $\phi$ admits an 1-in-3 assignment. Moreover, every matching cut of $G$, if any, is a perfect matching cut.
A general idea in constructing a graph without long induced paths is to ensure that long paths must pass through few cliques, say at most two. The following observation gives hints on how to create such ‘path-blocking’ cliques: consider a graph, in which the seven vertices $c,c_{k},a_{k}$, $1\leq k\leq 3$, induce a tree with leaves $a_{1},a_{2},a_{3}$ and degree-2 vertices $c_{1},c_{2},c_{3}$ and the degree-3 vertex $c$. If the graph has a perfect matching cut, then $a_{1},a_{2},a_{3}$ must belong to the same part of the cut. Therefore, we can make $\{a_{1},a_{2},a_{3}\}$ adjacent to a clique and the resulting graph still has a perfect matching cut. Now, a gadget $G(H;v)$ may be obtained from a suitable graph $H$ with $v\in V(H)$ as follows. Let $H$ be a graph having a degree-3 vertex $v$ with neighbors $b_{1},b_{2}$ and $b_{3}$. Let $G(H;v)$ be the graph obtained from $H-v$ by adding 7 new vertices $a_{1},a_{2},a_{3},\allowbreak c_{1},c_{2},c_{3}$ and $c$, and edges $cc_{k}$, $c_{k}a_{k}$, $a_{k}b_{k}$, $1\leq k\leq 3$, and $a_{1}a_{2},a_{1}a_{3}$ and $a_{2}a_{3}$. (Thus, contracting the triangle $a_{1}a_{2}a_{3}$ from $G(H;v)-\{c,c_{1},c_{2},c_{3}\}$ we obtain the graph $H$.) See Fig. 2.
@@
Assume that, for any neighbor $w$ of $v$ in $H$, $H$ has a perfect matching cut $(X,Y)$ such that $v\in X$ and $w\in Y$. Then, for any neighbor $d$ of $c$ in $G(H;v)$, the graph $G(H;v)$ has a perfect matching cut $(X^{\prime},Y^{\prime})$ such that $c\in X^{\prime}$ and $d\in Y^{\prime}$.
Examples of graphs $H$ in Observation 1 include the cube, the Petersen graph and the 10-vertex Heggernes-Telle graph in (HeggernesT98, , Fig. 3.1 with $v=a_{i}^{1}$). Our gadget will be obtained by taking the cube. [^2] Taking the Petersen graph or the Heggernes-Telle graph will produce induced $P_{t}$ for some $t\geq 15$. If there exists another graph $H$ ‘better’ than the cube, then our construction will yield a $P_{t}$-free graph for some $10\leq t\leq 13$.
@@
$T$$F$$F$$c_{1}$$x$$y$$z$$a_{11}$$a_{12}$$a_{13}$$b_{11}$$b_{12}$$b_{13}$$c_{11}^{\prime}$$c_{12}^{\prime}$$c_{13}^{\prime}$$c_{1}^{\prime}$$c_{2}$$u$$z$$y$$a_{21}$$a_{22}$$a_{23}$$b_{21}$$b_{22}$$b_{23}$$c_{21}^{\prime}$$c_{22}^{\prime}$$c_{23}^{\prime}$$c_{2}^{\prime}$$c_{3}$$z$$v$$w$$a_{31}$$a_{32}$$a_{33}$$b_{31}$$b_{32}$$b_{33}$$c_{31}^{\prime}$$c_{32}^{\prime}$$c_{33}^{\prime}$$c_{3}^{\prime}$
@@
Notice that no edge exists between the two ‘path-blocking’ cliques $F$ and $T$. Notice also that $G-F-T$ has exactly $m+n$ components:
  • For each $1\leq j\leq m$, the $6$-cycle $D_{j}:\,b_{j1},c_{j1}^{\prime},b_{j2},c_{j3}^{\prime},b_{j3},c_{j2}^{\prime}$ is a component of $G-F-T$, call it the clause $6$-cycle (of clause $C_{j}$);
  • For each variable $x$, the clique $Q(x)$ is a component of $G-F-T$, call it the variable clique (of variable $x$).
@@
First, observe that each component of $G-F-T$ is a clause $6$-cycle $D_{j}$ or a variable clique $Q(x)$. Hence,
$G-F-T$ is $P_{6}$-free. (1)
@@
Observe next that, for each $j$, $c_{j}^{\prime}\in F$ is the cut-vertex in $G-T$ separating the clause $6$-cycle $D_{j}$ and $F$, and $N(c_{j}^{\prime})\cap D_{j}=\{c_{j1}^{\prime},c_{j2}^{\prime},c_{j3}^{\prime}\}$. Observe also that, for each $x$, $(G-T)[Q(x)\cup F]$ is a co-bipartite graph, the complement of a bipartite graph. Hence, it can be verified immediately that
$G-T$ is $P_{7}$-free. (2)

### v3

@@ Abstract
In a graph, a (perfect) matching cut is an edge cut that is a (perfect) matching. matching cut (mc), respectively, perfect matching cut (pmc), is the problem of deciding whether a given graph has a matching cut, respectively, a perfect matching cut. The disconnected perfect matching problem (dpm) is to decide if a graph has a perfect matching that contains a matching cut. Solving an open problem posed in [Lucke, Paulusma, Ries (ISAAC 2022, Algorithmica 2023)], we show that pmc is $\mathsf{NP}$-complete in graphs without induced $14$-vertex path $P_{14}$. Our reduction also works simultaneously for mc and dpm, improving the previous hardness results of mc on $P_{15}$-free graphs and of dpm on $P_{19}$-free graphs to $P_{14}$-free graphs for both problems. Actually, we prove a slightly stronger result: within $P_{14}$-free $8$-chordal graphs (graphs without chordless cycles of length at least $9$), it is hard to distinguish between those without matching cuts (respectively, perfect matching cuts, disconnected perfect matchings) and those in which every matching cut is a perfect matching cut. Moreover, assuming the Exponential Time Hypothesis, none of these problems can be solved in $2^{o(n)}$ time for $n$-vertex $P_{14}$-free $8$-chordal graphs.
On the positive side, we show that, as for mc [Moshi (JGT 1989)], dpm and pmc are polynomially solvable when restricted to $4$-chordal graphs. Together with the negative results, this partly answers an open question on the complexity of pmc in $k$-chordal graphs asked in [Le, Telle (WG 2021, TCS 2022) & Lucke, Paulusma, Ries (MFCS 2023, TCS 2024)].
@@ 1 Introduction and results
In a graph $G=(V,E)$, a cut is a partition $V=X\cup Y$ of the vertex set into disjoint, non-empty sets $X$ and $Y$. The set of all edges in $G$ having an endvertex in $X$ and the other endvertex in $Y$, written $E(X,Y)$, is called the edge cut of the cut $(X,Y)$. A matching cut is an edge cut that is a (possibly empty) matching. Another way to define matching cuts is as follows; see [6, 12]: a cut $(X,Y)$ is a matching cut if and only if each vertex in $X$ has at most one neighbor in $Y$ and each vertex in $Y$ has at most one neighbor in $X$. The classical $\mathsf{NP}$-complete problem matching cut (mc[6] asks if a given graph admits a matching cut.
An interesting special case, where the edge cut $E(X,Y)$ is a perfect matching of $G$, was considered in [13]. Such a matching cut is called a perfect mathing cut and the perfect matching cut (pmc) problem asks whether a given graph admits a perfect matching cut. It was shown in [13] that this special case pmc of mc remains $\mathsf{NP}$-complete.
A notion related to matching cut is disconnected perfect matching which has been considered recently in [4]: a disconnected perfect matching is a perfect matching that contains a matching cut. Observe that any perfect matching cut is a disconnected perfect matching but not the converse. Fig. 1 provides some small examples for matching cuts, perfect matching cuts and disconnected perfect matchings.
@@
The related problem to mc and pmc, disconnected perfect matching (dpm), asks if a given graph has a disconnected perfect matching; equivalently: if a given graph has a matching cut that is extendable to a perfect matching. It was shown in [4] that dpm is $\mathsf{NP}$-complete. All these three problems have received much attention lately; see, e.g.,  [2, 4, 5, 8, 11, 19, 20, 21, 22, 23] for recent results. In particular, we refer to [22] for a short survey on known applications of matching cuts in other research areas.
In this paper, we focus on the complexity and algorithms of these three problems restricted to graphs without long induced paths and cycles. $P_{4}$-free graphs have already been considered by Bonsma [3]. He proved that the $4$-cycle is the only connected $P_{4}$-free graph with minimum degree at least two that has a matching cut. In [7], Feghali initiated the study of mc for graphs without a fixed path as an induced subgraph. He showed that mc is solvable in polynomial time for $P_{5}$-free graphs, but that there exists an integer $t>0$ for which it is $\mathsf{NP}$-complete for $P_{t}$-free graphs; indeed, $t=27$ as pointed out in [22]. The current best known hardness results for mc and dpm in graphs without long induced paths are:
Theorem 1 ([22])
@@
Prior to the present paper, no similar hardness result for pmc was known. Indeed, it was asked in [22], whether there is an integer $t$ such that pmc is $\mathsf{NP}$-complete in $P_{t}$-free graphs. Polynomial-time algorithms exist for mc, pmc and dpm in $P_{6}$-free graphs [22].
@@
Previously, no similar polynomial-time results for dpm and pmc in $4$-chordal were known. Indeed, the computational complexity of pmc in $k$-chordal graphs, $k\geq 4$, has been stated as an open problem in [19, 22].
Our contributions
We prove that pmc is $\mathsf{NP}$-complete in graphs without induced path $P_{14}$, solving the open problem posed in [22]. For mc and dpm  we improve the hardness results in Theorem 1 in graphs without induced path $P_{15}$, respectively, $P_{19}$, to graphs without induced path $P_{14}$. It is remarkable that all these hardness results for three problems will be obtained simultaneously by only one reduction. Thus, our approach unifies and improves known approaches and results in the literature. Our hardness results can be stated in more details as follows.
@@
Our polynomial-time algorithm solving dpm in $4$-chordal graphs is based on simple, known forcing rules. We remark that the approach also works for mc, providing another proof for Theorem 2. In case of pmc in $4$-chordal graphs, we will give a non-trivial polynomial-time reduction to 2sat, based on the distance-level structure of $4$-chordal graphs admitting perfect matching cuts. Our positive and negative results in Theorems 3 and 4 on $k$-chordal graphs partly answer an open question stated in [19, 22].
@@
$t$ $\leq 6$ 77 $\cdots$ 1313 $\geq 14$
mc $\mathsf{P}$ ? $\cdots$ ? $\mathsf{NPC}$$(\star)$
pmc $\mathsf{P}$ ? $\cdots$ ? $\mathsf{NPC}$$(\star)$
dpm $\mathsf{P}$ ? $\cdots$ ? $\mathsf{NPC}$$(\star)$
$k$ $\leq 4$ 55 66 77 $\geq 8$
mc $\mathsf{P}$ ? ? ? $\mathsf{NPC}$$(\star)$
pmc $\mathsf{P}$$(\star)$ ? ? ? $\mathsf{NPC}$$(\star)$
dpm $\mathsf{P}$$(\star)$ ? ? ? $\mathsf{NPC}$$(\star)$
Table 1: The current status of the computational complexity of matching cut problems in $P_{t}$-free graphs (left hand-side) and in $k$-chordal graphs (right hand-side). $\mathsf{P}$, $\mathsf{NPC}$, and ‘?’ means polynomial, $\mathsf{NP}$-complete, and unknown, respectively. Results of this paper are marked with $(\star)$.
@@ 2 Preliminaries
[…]
@@
If $G$ has a perfect matching cut $M=(X,Y)$ then $N[v]$ is not monochromatic for any vertex $v$. Indeed, every vertex $v\in X$ has exactly one neighbor in $Y$ which we sometimes call the private neighbor of $v$ and denote $p(v)$, and vice versa, every vertex $w\in Y$ has its private neighbor $p(w)$ in $X$. Thus, for any vertex $v$, $p(p(v))=v$.
Algorithmic lower bounds in this paper are conditional, based on the Exponential Time Hypothesis (ETH) [14]. The ETH asserts that no algorithm can solve 3sat in subexponential time $2^{o(n)}$ for $n$-variable 3-cnf formulas: there exists a constant $\delta>0$ such that no algorithm can solve 3sat in $2^{\delta n}$ time for 3-cnf formulas over $n$ variables. As shown by the Sparsification Lemma in [15], the hard cases of 3sat already consist of sparse formulas with $m=O(n)$ clauses. Hence, the ETH implies that 3sat cannot be solved in $2^{o(n+m)}$ time.
Recall that an instance for 1-in-3sat is a 3-cnf formula $\phi=C_{1}\land C_{2}\land\cdots\allowbreak\land C_{m}$ over $n$ variables, in which each clause $C_{j}$ consists of three distinct literals. The problem asks whether there is a truth assignment to the variables such that every clause in $\phi$ has exactly one true literal. We call such an assignment a 1-in-3 assignment. There is a polynomial reduction from 3sat to 1-in-3sat ([25, Theorem 7.2], see also [27]), which transforms an instance for 3sat to an equivalent instance for 1-in-3sat  linear in the number of variables and clauses. Thus, assuming ETH, 1-in-3sat cannot be solved in $2^{o(n+m)}$ time on inputs with $n$ variables and $m$ clauses. We will need a restriction of 1-in-3sat, positive 1-in-3sat, in which each variable occurs only positively. There is a well-known polynomial reduction from 1-in-3sat to positive 1-in-3sat, which transforms an instance for 1-in-3sat to an equivalent instance for positive 1-in-3sat, linear in the number of variables and clauses. (See also an exercise in [25, Exercise 7.1].) Hence, we obtain: assuming ETH, positive 1-in-3sat cannot be solved in $2^{o(n+m)}$ time for inputs with $n$ variables and $m$ clauses.
3 Hardness results: Proof of Theorem 3
Recall that a perfect matching cut is in particular a matching cut, as well as a disconnected perfect matching. This observation leads to the following promise versions of mc, pmc and dpm. (We refer to [9, 10] for background on promise problems.)
promise-pmc mc
Instance: A graph $G$ that either has no matching cut, or every matching cut is a perfect matching cut. Question: Does $G$ have a matching cut ?
promise-pmc pmc
Instance: A graph $G$ that either has no perfect matching cut, or every matching cut is a perfect matching cut. Question: Does $G$ have a perfect matching cut ?
promise-pmc dpm
Instance: A graph $G$ that either has no disconnected perfect matching, or every matching cut is a perfect matching cut. Question: Does $G$ have a disconnected perfect matching ?
@@ 3.1 The reduction
Let $\phi$ be a 3-cnf formula with $m$ clauses $C_{j}$, $1\leq j\leq m$, and $n$ variables $x_{i}$, $1\leq i\leq n$, in which each clause $C_{j}$ consists of three distinct variables. We will construct a $\{3P_{6},2P_{7},P_{14}\}$-free $8$-chordal graph $G$ such that $G$ has a perfect matching cut if and only if $\phi$ admits a 1-in-3 assignment. Moreover, every matching cut of $G$, if any, is a perfect matching cut.
A general idea in constructing a graph without long induced paths is to ensure that long paths must pass through few cliques, say at most two. The following observation gives hints on how to create such ‘path-blocking’ cliques: consider a graph, in which the seven vertices $c,c_{k},a_{k}$, $1\leq k\leq 3$, induce a tree with leaves $a_{1},a_{2},a_{3}$ and degree-2 vertices $c_{1},c_{2},c_{3}$ and the degree-3 vertex $c$. If the graph has a perfect matching cut, then $a_{1},a_{2},a_{3}$ must belong to the same part of the cut. Therefore, we can make $\{a_{1},a_{2},a_{3}\}$ adjacent to a clique and the resulting graph still has a perfect matching cut. Now, a gadget $G(H;v)$ may be obtained from a suitable graph $H$ with $v\in V(H)$ as follows. Let $H$ be a graph having a degree-3 vertex $v$ with neighbors $b_{1},b_{2}$ and $b_{3}$. Let $G(H;v)$ be the graph obtained from HvH-v by adding 7 new vertices $a_{1},a_{2},a_{3},\allowbreak c_{1},c_{2},c_{3}$ and $c$, and edges $cc_{k}$, $c_{k}a_{k}$, $a_{k}b_{k}$, $1\leq k\leq 3$, and $a_{1}a_{2},a_{1}a_{3}$ and $a_{2}a_{3}$. (Thus, contracting the triangle $a_{1}a_{2}a_{3}$ from $G(H;v)-\{c,c_{1},c_{2},c_{3}\}$ we obtain the graph $H$.) See Fig. 2.
[…]
@@
Assume that, for any neighbor $b_{i}$ of $v$ in $H$, $H$ has a perfect matching cut $(X,Y)$ such that $v\in X$ and $b_{i}\in Y$. Then, for any neighbor $d$ of $c$ in $G(H;v)$, the graph $G(H;v)$ has a perfect matching cut $(X^{\prime},Y^{\prime})$ such that $c\in X^{\prime}$ and $d\in Y^{\prime}$.
Examples of graphs $H$ in Observation 1 include the cube, the Petersen graph and the 10-vertex Heggernes-Telle graph in [13, Fig. 3.1 with $v=a_{i}^{1}$]. Our gadget will be obtained by taking the cube. [^2] Taking the Petersen graph or the Heggernes-Telle graph will produce induced $P_{t}$ for some $t\geq 15$. If there exists another graph $H$ ‘better’ than the cube, then our construction will yield a $P_{t}$-free graph for some $10\leq t\leq 13$.
@@
[…]
@@
Notice that no edge exists between the two ‘path-blocking’ cliques $F$ and $T$. Notice also that GFTG-F-T has exactly $m+n$ components:
  • For each $1\leq j\leq m$, the $6$-cycle $D_{j}:\,b_{j1},c_{j1}^{\prime},b_{j2},c_{j3}^{\prime},b_{j3},c_{j2}^{\prime}$ is a component of GFTG-F-T, call it the clause $6$-cycle (of clause $C_{j}$);
  • For each variable $x$, the clique $Q(x)$ is a component of GFTG-F-T, call it the variable clique (of variable $x$).
@@
First, observe that each component of GFTG-F-T is a clause $6$-cycle $D_{j}$ or a variable clique $Q(x)$. Hence,
GFTG-F-T is $P_{6}$-free. (1)
@@
Observe next that, for each $j$, $c_{j}^{\prime}\in F$ is the cut-vertex in GTG-T separating the clause $6$-cycle $D_{j}$ and $F$, and $N(c_{j}^{\prime})\cap D_{j}=\{c_{j1}^{\prime},c_{j2}^{\prime},c_{j3}^{\prime}\}$. Observe also that, for each $x$, $(G-T)[Q(x)\cup F]$ is a co-bipartite graph, the complement of a bipartite graph. Hence, it can be verified immediately that
GTG-T is $P_{7}$-free. (2)